OpenStudy (anonymous):

Solve. 4^(x + 7) = 6^(x – 1) x ≈ 0.04 x ≈ 0.05 x ≈ 19.51 x ≈ 28.35

OpenStudy (anonymous):

take natural log of both sides and solve for x

OpenStudy (anonymous):

(x+7) ln 4 = (x-1) ln 6 x ln 4 + 7 ln 4 = x ln 6 - 1 ln 6 x ln 4 -x ln 6 = -7 ln 4 -1 ln 6 x (ln 4 -ln 6) = -11.62 x = -11.62/-.41 x = 28.35